### v1

@@ Abstract
Recently, therapeutic peptides have demonstrated great promise for cancer treatment. To explore powerful anticancer peptides, artificial intelligence (AI)-based approaches have been developed to systematically screen potential candidates. However, the lack of efficient featurization of peptides has become a bottleneck for these machine-learning models. In this paper, we propose a topology-enhanced machine learning model (Top-ML) for anticancer peptide prediction. Our Top-ML employs peptide topological features derived from its sequence "connection" information characterized by vector and spectral descriptors. Our Top-ML model has been validated on two widely used AntiCP 2.0 benchmark datasets and has achieved state-of-the-art performance. Our results highlight the potential of leveraging novel topology-based featurization to accelerate the identification of anticancer peptides.
Availability: The Top-ML code can be found in https://github.com/XueGong-git/TopPep-ML. The datasets for ACPs can be found in https://webs.iiitd.edu.in/raghava/anticp2/download.php.
Contact: weejunji@msu.edu, xue.gong@ntu.edu.sg, xiakelin@ntu.edu.sg
@@ Introduction
Despite the extensive research into the prevention and diagnosis of cancer over the past few decades, cancer remains one of the leading causes of death worldwide. Statistics from the World Health Organization [1] show that cancer accounted for an estimated 10 million deaths in 2020, representing nearly one in six deaths. Traditional approaches to cancer treatment often come with significant limitations, such as systemic toxicity and drug resistance. Traditional cancer treatment relies on radiotherapy and chemotherapy; however, both methods have downsides, as they target both normal and cancerous cells. Furthermore, highly adaptable cancer cells can develop drug resistance, further reducing the efficacy of these treatments.
@@
In recent years, there has been a growing interest in Anticancer Peptides (ACPs) as a novel alternative treatment for cancer. ACPs are a specific type of antimicrobial peptides (AMP) typically consisting of less than 50 amino acids. They differ from other AMPs by their cationic nature and low molecular weight [2]. Peptide therapy offers significant advantages over traditional cancer treatments due to its high specificity, low toxicity, superior membrane penetration ability, and easier chemical modifications. However, the experiment-based discovery and design of ACPs are extremely costly, time-consuming, and labor-intensive, which leads to poor scalability. To address these challenges, researchers are increasingly turning to machine learning for the screening and design of ACPs.
@@
Featurization or feature engineering is key to the performance of machine learning models for ACPs design and discovery. The ACPs identification task was formulated as a binary classification machine learning problem in these studies [3, 4, 5, 6, 7, 8]. In these methods, features are constructed from the peptides and then used to train ML models. Traditional molecular descriptors are primarily designed based on molecular structure properties, such as amino acid composition, dipeptide composition, and terminus composition to establish strong structure-function relationships [5]. In addition, other classes of physicochemical properties, including hydrophobicity and amphipathicity, have been utilized as feature encoding schemes [9]. The effectiveness of machine learning techniques heavily depends on molecular featurization and representations [10, 11, 12].
@@
Recently, topological data analysis (TDA) [13, 14], which is a model deeply rooted in algebraic topology, has demonstrated its great power in molecular representation and featurization [15]. In particular, TDA-based machine learning models have achieved great successes in various steps of drug design, including protein-ligand binding affinity prediction [15], protein stability change upon mutation prediction [16], toxicity prediction [17], solvation free energy prediction [18], partition coefficient and aqueous solubility [19], binding pocket detection [20], protein mutation analysis [21], and drug discovery [22]. In the D3R Grand Challenge, a worldwide competition for computational models in drug design, TDA-based models have consistently delivered state-of-the-art results [23]. Further, TDA-based models have also achieved great successes in material data analysis [24, 25]. The key idea of these models is to characterize the structural "connection" information with topological tools.
@@
Here, we propose a topology-enhanced machine learning model (Top-ML) for ACPs prediction. The key idea is to use topological features to characterize the ‘‘connection’’ pattern between amino acids within peptide sequences. More specifically, we employ vector features and spectral features derived from amino acid sequences to describe these connection properties, i.e., how amino acids, amino acid pairs, and amino acid groups relate to each other. Our Top-ML considers four types of peptide features, namely Magnus vector, natural vector, terminal composition feature, and spectral feature. By combining all four types of features with the Extra Trees classifier, we demonstrate that our proposed Top-ML can achieve state-of-the-art results in the widely-used benchmark datasets.
@@ Results
Natural vector has been widely used in the construction of phylogenetic trees [26, 27, 28]. Here we use natural vector representation for peptide sequences. The natural vector was first conceived for the analysis of genomes and has been previously applied to DNA sequences [29]. Consider a peptide sequence $\mathcal{S}=x_{1}x_{2}\cdots x_{n}$ of length $n$, where $x_{i}$ lies in the set of 20 unique amino acids $\{A,C,D,\cdots,Y\}$. For each peptide sequence $\mathcal{S}$ and amino acid $R$, we compute three statistical metrics: $n_{R}$, representing the number of times $R$ appears in $\mathcal{S}$; $\mu_{R}$, the average position of $R$; and $D_{2}^{R}$, the variation of position. The natural vector concatenates the three statistics calculated for each unique amino acid, forming a vector of length 60, i.e., ($n_{A}$, $\mu_{A}$, $D_{2}^{A}$, … $n_{R}$, $\mu_{R}$, $D_{2}^{R}$, … $n_{Y}$, $\mu_{Y}$, and $D_{2}^{Y}$). The natural vector encodes the presence and location of amino acids which correlate with anticancer properties. For instance, a previous study found that A, F, H, K, L, and W are more prevalent in ACPs, whereas C, G, R, and S are more common in non-ACPs [5].
@@
In combinatorial group theory, Wilhelm Magnus studied representations of free groups by non-commutative power series. For a free group $F$ with basis $x_{1},\cdots,x_{n}$ and a power series ring $\prod$ in indeterminates $\xi_{1},\cdots,\xi_{n}$, Magnus proved that the map $\mu:\xi_{i}\mapsto 1+\xi_{i}$ defines an isomorphism from $F$ into the multiplicative group $\prod^{\times}$ of units in $\prod$ [30]. Magnus’s work laid the foundation for the development of Magnus representation, which has been recently applied to the analysis of DNA and RNA sequences in genome analysis [31].
@@
In implementation, the Magnus vector of a peptide sequence $S$, denoted as $v(S)$, is obtained by the following steps. First, we arrange the set of possible words over the amino acids $\{A,C,D,\cdots,W,Y\}$ of length less than or equal to $n$, first by ascending order of length and then by lexicographic order. With respect to the above arrangement, assign each term present in $\rho(S)$ with a constant coefficient $c\in R$, and 0 for each term not present in $\rho(S)$. In our model, we use $c=1$ and only consider subsequences with a maximum length of 2. As each peptide sequence can contain a maximum of 20 unique amino acids [5], there are a maximum of $\sum_{i=1}^{2}20^{i}=420$ possible subsequences of up to length 2, i.e., $A,C,D,\cdots,W,Y,AA,AC,AD,\cdots,YW,YY$. Note that due to the non-commutativity of the variables, if $S^{\prime}=ADC$, we observe that $\rho(S^{\prime})\neq\rho(S)$ and $v(S^{\prime})\neq v(S)$. From our numerical experiments detailed in the Results section, we found that the Magnus vector of length 420 are sufficient to achieve state-of-the-art results in anticancer peptide classification.
Figure 1: Illustration of the transformation of (a) the peptide sequence ‘‘ACDE’’ into (b) a sequence-based Laplacian $\mathbf{L}_{0}$ and (c) its corresponding symmetric normalized Laplacian matrix $\tilde{\mathbf{L}}_{0}$ based on the mean index positions of the amino acids. The largest off-diagonal entry in (b) $\mathbf{L}_{0}$ in terms of absolute value corresponds to the amino acid pair (D, E) since D and E have the highest mean index positions. After normalization, the diagonal entries in (c) $\tilde{\mathbf{L}}_{0}$ become ones, and the off-diagonal entries take values between -1 and 0.
In our implementation, we consider the Magnus vectors for non-overlapping $k$-mers for each peptide sequence. A $k$-mer represents a segment of $k$ consecutive amino acids within a given peptide sequence. We compute the Magnus vector for each $k$-mer, selected using non-overlapping sliding windows of size $k$ to avoid redundancy and reduce computational cost. The sliding window begins at the first index, and we truncate the tail if the remaining sequence is shorter than $k$. Therefore, the number of $k$-mers for a peptide sequence of length $n$ is the largest integer less than or equal to $\frac{n}{k}$. We then form a mean Magnus vector $m(S)$ by averaging the Magnus vectors across all $k$-mers, capturing both $k$-mer frequency information and internal subsequence structure. We experiment with sliding window sizes $k=4,5,6,7$ and find that $k=5$ yields good prediction performance for AntiCP 2.0 Datasets, as detailed in the Supporting Information Table 3. This result aligns with the observation of minimal classification errors for the Baltimore and genus classification labels reported in [32]. Therefore, we selected $k=5$ for our final Top-ML model.
@@
Previous research into peptide representation has found that terminal residues provide vital information about the biological and physiological functions of peptides [33]. The terminus composition provides local sequence information in contrast to the global sequential information when the full length of the peptide is used for featurization. We compute the Magnus and natural vectors based on the terminus compositions of peptides (i.e., N5, C5, N5C5, $\cdots$, C15, N15C15). For example, extracting the first 5 residues from the N-terminus is termed N5. Our experiments show that the C15 and N15C15 terminus provided the best prediction accuracy for AntiCP 2.0 Datasets A and B respectively, as shown in Table 5 in the Supplementary Information.
@@
In this section, we introduce a peptide spectral representation, which is inspired by the spectral theory of combinatorial Laplacian matrices. Simplicial complexes offer an effective representation for capturing the underlying structural information and its interactions. We refer the reader to the Method section for more information regarding simplicial complexes and their combinatorial Laplacian matrices. Using simplicial complexes as a molecular structural representation has generated tremendous success for combinatorial Laplacian matrices in applications such as drug design and discovery [34]. However, unlike molecular structures, which inherently possess 3D structures, peptide data is sequential and cannot be directly represented by various topological objects such as graphs and simplicial complexes. Therefore, we need to first construct simplicial complexes from peptide sequences.
@@
In order to construct a peptide spectral representation, we first define a sequence-based boundary matrix. For any peptide sequence $S$ of length $n$, we denote the set of possible subsequences of length $k$ as $S_{k}$, where $1\leq k\leq n$. Then the sequence-based boundary matrix $\mathbf{B}_{k}$ ($1\leq k\leq n$) is an $|S_{k}|\times|S_{k+1}|$ matrix constructed from the relationships between subsequences $w^{k}\in S_{k}$ and $w^{k+1}\in S_{k+1}$. For a peptide sequence $S$ of length $n$, its $k$-th sequence-based boundary matrix $\mathbf{B}_{k}$ is defined as follows:
$\mathbf{B}_{k}(i,j)=\begin{cases}f(w_{i}^{k}),&\text{if }w_{i}^{k}\preceq w_{j}^{k+1},\\ 0,&\text{if }w_{i}^{k}\not\preceq w_{j}^{k+1},\end{cases}$ (1)
@@
where $f(w)$ is a real-valued function, $w_{i}^{k}\preceq w_{j}^{k+1}$ means that the $w_{i}^{k}$ is a subsequence of $w_{j}^{k+1}$. Likewise, $w_{i}^{k}\not\preceq w_{j}^{k+1}$ means that $w_{i}^{k}$ is not a subsequence of $w_{j}^{k+1}$. We are considering two definitions for $f(w)$. The first defines $f(w)$ as the number of occurrences of the subsequence $w$ within the sequence $S$. For instance, in the sequence ‘‘FFSFS", the subsequence ‘‘FS" occurs twice. Another definition of $f(w)$ is the mean index position of the first amino acid in $w$ across all its occurrences within sequence $S$. For example, in ‘‘FFSFS", the index positions of ‘‘FS" are 2 and 4, resulting in a mean index position of 3. The choice of $f(w)$ for the final model is the mean index position based on the model performance as presented in the Supplementary Information Table 2.
@@
Similar to the combinatorial Laplacian defined on simplicial complexes, we compute our sequence-based Laplacian matrices $\mathbf{L}_{k}$ using Equation (3) in the Method section. For $\mathbf{L}_{k}$ to be positive semi-definite, we update the diagonals of $\mathbf{L}_{k}$ using the sum of its off-diagonals and then assign all off-diagonal values a negative sign. This ensures the sum of every row in $\mathbf{L}_{k}$ is zero. We then convert $\mathbf{L}_{k}$ to its symmetric normalized version $\tilde{\mathbf{L}}_{k}$ as defined in Equation (4).
@@
Let’s consider $\mathbf{B}_{1}$ as an example. Given that all peptides consist of a maximum of 20 unique amino acids [5], the rows in $\mathbf{B}_{1}$ represent these 20 amino acids (A, C, D, …, Y), while the columns represent the 400 pairs of amino acids (AA, AC, AD, …, YY). Therefore, $\mathbf{B}_{1}$ is a $20\times 400$ matrix, $\mathbf{L}_{0}$ is a $20\times 20$ matrix, and $\mathbf{L}_{1}$ is a $400\times 400$ matrix. This resulting vector containing eigenvalues is used as input features for downstream machine learning models. $\mathbf{L}_{0}$ produces a feature vector of size 20 while $\mathbf{L}_{1}$ produces a feature vector of size 400. The two feature vectors are concatenated to form a feature vector of size 420.
@@
In this section, we present our topology-enhanced machine learning model (Top-ML) and its results for anticancer peptide classification. The natural vector, Magnus vector, terminus composition features, and peptide spectral representations are combined to form the input features for our Top-ML model. Figure 2 shows the general pipeline of our Top-ML model.
@@
We train and assess the performance of our Top-ML model using the AntiCP 2.0 benchmark datasets [5]. The AntiCP 2.0 datasets comprise Dataset A and Dataset B. Dataset A consists of 970 experimentally validated ACPs and 970 randomly selected peptides, assumed to be non-ACPs. Dataset B contains 861 experimentally validated ACPs and 861 non-ACPs, which are antimicrobial peptides (AMPs). AntiCP 2.0 (Datasets A and B) was partitioned into training and test sets using an 80/20 split. We assess the effectiveness of our Top-ML model by employing a range of performance metrics, which include accuracy, sensitivity (true positive rate), specificity (true negative rate), and the Matthews correlation coefficient (MCC). The model selection process is detailed in the Supplementary Information. The final Top-ML model for AntiCP 2.0 is an Extra Trees classifier with 400 trees, each using a random subset of features equal to the square root of the total number of features to determine the optimal split. In this finalized model, peptide spectral representations are derived from the mean index position, and Magnus vectors are computed using a window size of 5. The terminus used to build terminal composition features is C15 for AntiCP 2.0 (Dataset A) and N15C15 for AntiCP 2.0 (Dataset B). The final Top-ML model has been further optimized by finding the optimal probability threshold that returns the best accuracy. All reported performances of the Top-ML model are medians over 100 repeated tests.
@@
Figure 2: Pipeline of topology-enhanced machine learning model (Top-ML) for anticancer peptide prediction. (A) The peptide sequences from AntiCP 2.0 (Dataset A or B). (B) The peptide spectral features are obtained from the sequence-based Laplacian matrices. (C) The Magnus vectors, natural vectors, and terminal composition features are generated from peptide sequences. (D) The combined features serve as inputs for the classification of anticancer peptides (ACPs) and non-ACPs.
@@
Table 1 summarizes the results of Top-ML and existing state-of-the-art machine learning models on AntiCP 2.0 (Datasets A and B). It can be observed that on AntiCP 2.0 (Dataset A), our Top-ML model achieves one of the highest accuracies and MCC scores, outperforming several existing state-of-the-art machine learning models. This is noteworthy because these models were built on deep learning algorithms while only tree-based classifiers have been used for Top-ML. Compared to AntiCP 2.0 [5], which serves as a key benchmark model for comparison, our Top-ML model demonstrates an improvement of 1.3% in accuracy and 3.1% in specificity. On AntiCP 2.0 (Dataset B), the accuracy of Top-ML stands at 78.89%, which is higher than all other models compared, except for iACP-FSCM (82.50% accuracy) and ME-ACP(79.2%). When compared with the benchmark model AntiCP 2.0, there is an improvement of 3.5% in accuracy and 8.5% in specificity. Although iACP-FSCM and ME-ACP report higher accuracy, our Top-ML model can obtain a more balanced sensitivity and specificity score.

### v4

@@ Abstract
Recently, therapeutic peptides have demonstrated great promise for cancer treatment. To explore powerful anticancer peptides, artificial intelligence (AI)-based approaches have been developed to systematically screen potential candidates. However, the lack of efficient featurization of peptides has become a bottleneck for these machine-learning models. In this paper, we propose a topology-enhanced machine learning model (Top-ML) for anticancer peptides prediction. Our Top-ML employs peptide topological features derived from its sequence “connection” information characterized by vector and spectral descriptors. Our Top-ML model, employing an Extra-Trees classifier, has been validated on the AntiCP 2.0 and mACPpred 2.0 benchmark datasets, achieving state-of-the-art performance or results comparable to existing deep learning models, while providing greater interpretability. Our results highlight the potential of leveraging novel topology-based featurization to accelerate the identification of anticancer peptides.
keywords:
American Chemical Society,
[…]
@@ Introduction
Despite the extensive research into the prevention and diagnosis of cancer over the past few decades, cancer remains one of the leading causes of death worldwide. Statistics from the World Health Organization 1 show that cancer accounted for an estimated 10 million deaths in 2020, representing nearly one in six deaths. Traditional approaches to cancer treatment often come with significant limitations, such as systemic toxicity and drug resistance. Traditional cancer treatment relies on radiotherapy and chemotherapy; however, both methods have downsides, as they target both normal and cancerous cells. Furthermore, highly adaptable cancer cells can develop drug resistance, further reducing the efficacy of these treatments.
@@
In recent years, there has been a growing interest in anticancer peptides (ACPs) as a novel alternative treatment for cancer. ACPs are a specific type of antimicrobial peptides (AMP) typically consisting of less than 50 amino acids. They differ from other AMPs by their cationic nature and low molecular weight 2. Peptide therapy offers significant advantages over traditional cancer treatments due to its high specificity, low toxicity, superior membrane penetration ability, and easier chemical modifications. However, the experiment-based discovery and design of ACPs are extremely costly, time-consuming, and labor-intensive, which leads to poor scalability. To address these challenges, researchers are increasingly turning to machine learning for the screening and design of ACPs.
@@
Featurization or feature engineering is key to the performance of machine learning models for ACPs design and discovery. The ACPs identification task was formulated as a binary classification machine learning problem in these studies 3, 4, 5, 6, 7, 8. In these methods, features are constructed from the peptides and then used to train ML models. Traditional molecular descriptors are primarily designed based on sequence composition, such as amino acid composition, dipeptide composition, and terminus composition to establish strong structure-function relationships 5. In addition, other classes of physicochemical properties, including hydrophobicity and amphipathicity, have been utilized as feature encoding schemes 9. The effectiveness of machine learning techniques heavily depends on molecular featurization and representations 10, 11, 12.
@@
Recently, topological data analysis (TDA) 13, 14, which is a model deeply rooted in algebraic topology, has demonstrated its great power in molecular representation and featurization 15. In particular, TDA-based machine learning models have achieved great successes in various steps of drug design, including protein-ligand binding affinity prediction 15, protein stability change upon mutation prediction 16, toxicity prediction 17, solvation free energy prediction 18, partition coefficient and aqueous solubility 19, binding pocket detection 20, protein mutation analysis 21, and drug discovery 22. In the D3R Grand Challenge, a worldwide competition for computational models in drug design, TDA-based models have consistently delivered state-of-the-art results 23. Further, TDA-based models have also achieved great successes in material data analysis 24, 25. The key idea of these models is to characterize the structural “connection” information with topological tools.
@@
Here, we propose a topology-enhanced machine learning model (Top-ML) for ACPs prediction. The key idea is to use topological features to characterize the “connection” pattern between amino acids within peptide sequences. More specifically, we employ vector features and spectral features derived from amino acid sequences to describe these connection properties, i.e., how amino acids, amino acid pairs, and amino acid groups relate to each other. Our Top-ML considers four types of peptide features, namely natural vector 26, Magnus vector 27, terminal composition feature, and spectral feature. By combining all four feature types with an Extra-Trees classifier, we demonstrate that the proposed Top-ML achieves state-of-the-art results on the alternative dataset of AntiCP 2.0 and accuracy comparable to the best deep learning models on the main dataset of AntiCP 2.0 and mACPpred 2.0. Additionally, our model offers greater interpretability, which we demonstrate through the feature importance analysis shown in Figure 3 using the mACPpred 2.0 dataset.
@@ Results
Natural vector has been widely used in the construction of phylogenetic trees 28, 29, 30. Here we use natural vector representation for peptide sequences. The natural vector was first conceived for the analysis of genomes and has been previously applied to DNA sequences 31. Consider a peptide sequence $\mathcal{S}=x_{1}x_{2}\cdots x_{n}$ of length $n$, where $x_{i}$ lies in the set of 20 unique amino acids $\{A,C,D,\cdots,Y\}$. For each peptide sequence $\mathcal{S}$ and amino acid $R$, we compute three statistical metrics: $n_{R}$, representing the number of times $R$ appears in $\mathcal{S}$; $\mu_{R}$, the average position of $R$; and $D_{2}^{R}$, the variation of position. The natural vector concatenates the three statistics calculated for each unique amino acid, forming a vector of length 60, i.e., ($n_{A}$, $\mu_{A}$, $D_{2}^{A}$, … $n_{R}$, $\mu_{R}$, $D_{2}^{R}$, … $n_{Y}$, $\mu_{Y}$, and $D_{2}^{Y}$). The natural vector encodes the presence and location of amino acids which correlate with anticancer properties. For instance, a previous study found that A, F, H, K, L, and W are more prevalent in ACPs, whereas C, G, R, and S are more common in non-ACPs 5.
@@
In combinatorial group theory, Wilhelm Magnus studied representations of free groups by non-commutative power series. For a free group $F$ with basis $x_{1},\cdots,x_{n}$ and a power series ring $\prod$ in indeterminates $\xi_{1},\cdots,\xi_{n}$, Magnus proved that the map $\mu:\xi_{i}\mapsto 1+\xi_{i}$ defines an isomorphism from $F$ into the multiplicative group $\prod^{\times}$ of units in $\prod$ 32. Magnus’s work laid the foundation for the development of Magnus representation, which has been recently applied to the analysis of DNA and RNA sequences in genome analysis 27.
@@
In implementation, the Magnus vector of a peptide sequence $S$, denoted as $v(S)$, is obtained by the following steps. First, we arrange the set of possible words over the amino acids $\{A,C,D,\cdots,W,Y\}$ of length less than or equal to $n$, first by ascending order of length and then by lexicographic order. With respect to the above arrangement, assign each term present in $\rho(S)$ with a constant coefficient $c\in R$, and 0 for each term not present in $\rho(S)$. In our model, we use $c=1$ and only consider subsequences with a maximum length of 2. As each peptide sequence can contain a maximum of 20 unique amino acids 5, there are a maximum of $\sum_{i=1}^{2}20^{i}=420$ possible subsequences of up to length 2, i.e., $A,C,D,\cdots,W,Y,AA,AC,AD,\cdots,YW,YY$.
Consider the peptide sequence $\mathcal{S}=ACD$. The subsequences present in $\mathcal{S}$ are $A,C,D,AC,$ and $CD$. The corresponding Magnus vector $v(S)=(1,1,1,0,...,0)^{T}$ is a 420-dimensional binary vector with 1s at the positions corresponding to these 5 subsequences and 0s elsewhere. Note that due to the non-commutativity of the variables, if $S^{\prime}=ADC$, we observe that $\rho(S^{\prime})\neq\rho(S)$ and $v(S^{\prime})\neq v(S)$. From our numerical experiments detailed in the Results section, we found that the Magnus vector of length 420 are sufficient to achieve state-of-the-art results in ACPs classification.
In our implementation, we consider the Magnus vectors for non-overlapping $k$-mers for each peptide sequence. A $k$-mer represents a segment of $k$ consecutive amino acids within a given peptide sequence. We compute the Magnus vector for each $k$-mer, selected using non-overlapping sliding windows of size $k$ to avoid redundancy and reduce computational cost. The sliding window begins at the first index, and we truncate the tail if the remaining sequence is shorter than $k$. Therefore, the number of $k$-mers for a peptide sequence of length $n$ is the largest integer less than or equal to $\frac{n}{k}$. We then form a mean Magnus vector $m(S)$ by averaging the Magnus vectors across all $k$-mers, capturing both $k$-mer frequency information and internal subsequence structure. We experiment with sliding window sizes $k=4,5,6,7$ and find that $k=5$ yields good prediction performance for AntiCP 2.0 Datasets, as detailed in Table 4 in the Methods section. This result aligns with the observation of minimal classification errors for the Baltimore and genus classification labels reported in 33. Therefore, we selected $k=5$ for our final Top-ML model.
@@
Previous research into peptide representation has found that terminal residues provide vital information about the biological and physiological functions of peptides 34. The terminus composition provides local sequence information in contrast to the global sequential information when the full length of the peptide is used for featurization. We compute the Magnus and natural vectors based on the terminus compositions of peptides (i.e., N5, C5, N5C5, $\cdots$, C15, N15C15). For example, extracting the first 5 residues from the N-terminus is termed N5. Our experiments show that the C15 and N15C15 terminus provided the best prediction accuracy for AntiCP 2.0 Datasets A and B respectively, as shown in Table 6.
@@
In this section, we introduce a peptide spectral representation, which is inspired by the spectral theory of combinatorial Laplacian matrices. Simplicial complexes offer an effective representation for capturing the underlying structural information and its interactions. We refer the reader to the Method section for more information regarding simplicial complexes and their combinatorial Laplacian matrices. Using simplicial complexes as a molecular structural representation has generated tremendous success for combinatorial Laplacian matrices in applications such as drug design and discovery 35. However, unlike molecular structures, which inherently possess 3D structures, peptide data is sequential and cannot be directly represented by various topological objects such as graphs and simplicial complexes. Therefore, we need to first construct simplicial complexes from peptide sequences.
@@
In order to construct a peptide spectral representation, we first define a sequence-based boundary matrix. For any peptide sequence $S$ of length $n$, we denote the set of possible subsequences of length $k$ as $S_{k}$, where $1\leq k\leq n$. Then the sequence-based boundary matrix $\hat{\mathbf{B}}_{k}$ ($1\leq k\leq n$) is an $|S_{k}|\times|S_{k+1}|$ matrix constructed from the relationships between subsequences $w^{k}\in S_{k}$ and $w^{k+1}\in S_{k+1}$. For a peptide sequence $S$ of length $n$, its $k$-th sequence-based boundary matrix $\hat{\mathbf{B}}_{k}$ is defined as follows:
$\hat{\mathbf{B}}_{k}(i,j)=\begin{cases}f(w_{i}^{k}),&\text{if }w_{i}^{k}\preceq w_{j}^{k+1},\\ 0&\text{otherwise },\end{cases}$ (1)
where $f(w)$ is a real-valued function, $w_{i}^{k}\preceq w_{j}^{k+1}$ means that the $w_{i}^{k}$ is a subsequence of $w_{j}^{k+1}$. The simplest choice is setting $f(w)=1$, in which case we obtain the standard boundary matrix in Equation (1), where entry $\hat{\mathbf{B}}_{k}(i,j)$ is 1 if $w_{i}^{k}$ appear as a subsequence of $w_{j}^{k+1}$ for at least once, and 0 otherwise. However, since a peptide sequence consists of 20 possible amino acid types, this would result in many different sequences having identical boundary matrices. To enhance expressive power and better differentiate peptide sequences, we incorporate both frequency and positional information by considering two definitions for $f(w)$.
@@
The first defines $f(w)$ as the number of occurrences of the subsequence $w$ within the sequence $S$. For instance, in the sequence “FFSFS”, the subsequence “FS” occurs twice. Another definition of $f(w)$ is the mean index position of the first amino acid in $w$ across all its occurrences within sequence $S$. For example, in “FFSFS”, the index positions of “FS” are 2 and 4, resulting in a mean index position of 3. Figure 1 shows another example for the construction of the sequence-based Laplacian using the mean index position. The choice of $f(w)$ for the final model is the mean index position based on the model performance as presented in Table 3.
Figure 1: Illustration of the transformation of (a) the peptide sequence “ACDE” into (b) a sequence-based Laplacian $\hat{\mathbf{L}}_{0}$ and (c) its corresponding symmetric normalized Laplacian matrix $\tilde{\mathbf{L}}_{0}$ based on the mean index positions of the amino acids. The largest off-diagonal entry in (b) $\hat{\mathbf{L}}_{0}$ in terms of absolute value corresponds to the amino acid pair (D, E) since D and E have the highest mean index positions. After normalization, the diagonal entries in (c) $\tilde{\mathbf{L}}_{0}$ become ones, and the off-diagonal entries take values between -1 and 0.
Similar to the combinatorial Laplacian defined on simplicial complexes, we compute our sequence-based Laplacian matrices $\hat{\mathbf{L}}_{k}$ using Equation (2) below:
$\hat{\mathbf{L}}_{k}=\begin{cases}\hat{\mathbf{B}}_{1}\hat{\mathbf{B}}_{1}^{T},&\text{if }k=0\\ \hat{\mathbf{B}}_{k}^{T}\hat{\mathbf{B}}_{k}+\hat{\mathbf{B}}_{k+1}\hat{\mathbf{B}}_{k+1}^{T},&\text{if }k\geq 1.\end{cases}$ (2)
@@
For $\hat{\mathbf{L}}_{k}$ to be positive semi-definite, we update the diagonals of $\hat{\mathbf{L}}_{k}$ using the sum of its off-diagonals and then assign all off-diagonal values a negative sign. This ensures the sum of every row in $\hat{\mathbf{L}}_{k}$ is zero. We then convert $\hat{\mathbf{L}}_{k}$ to its symmetric normalized version $\tilde{\mathbf{L}}_{k}$ as defined in Equation (5).
@@
Let’s consider $\hat{\mathbf{B}}_{1}$ as an example. Given that all peptides consist of a maximum of 20 unique amino acids 5, the rows in $\hat{\mathbf{B}}_{1}$ represent these 20 amino acids (A, C, D, …, Y), while the columns represent the 400 pairs of amino acids (AA, AC, AD, …, YY). Therefore, $\tilde{\mathbf{L}}_{1}$ is a $20\times 400$ matrix, $\tilde{\mathbf{L}}_{0}$ is a $20\times 20$ matrix, and $\tilde{\mathbf{L}}_{1}$ is a $400\times 400$ matrix. This resulting vector containing eigenvalues is used as input features for downstream machine learning models. $\tilde{\mathbf{L}}_{0}$ produces a feature vector of size 20 while $\tilde{\mathbf{L}}_{1}$ produces a feature vector of size 400. The two feature vectors are concatenated to form a feature vector of size 420.
ur Laplacian can be viewed as a weighted Laplacian matrix, where the weights reflect either the frequency or position at which subsequence pairs appear together. $\tilde{\mathbf{L}}_{0}$ resembles a graph Laplacian where nodes represent amino acids, and edges indicate consecutive appearances. Similarly, in $\tilde{\mathbf{L}}_{1}$, nodes represent dipeptides (or 2-mers), with edges connecting dipeptides that are adjacent and overlap by one amino acid. Such $k$-mers representation are widely used in bioinformatics to cluster amino acid sequences 36. The degeneracy of the zero eigenvalue counts the number of disconnected components of the underlying graph, while a smaller $p$-th eigenvalue indicates better $p$-way clustering 37. In our setting, good clustering means that subsequences within the same cluster frequently appear together, while those from different clusters are more separated. Biologically, such clustering may reflect amino acid groupings based on physicochemical properties, such as polar, neutral, or hydrophobic. Previously, adjacency pattern between these groups (referred to as Transition) have been used to predict protein folding 38 and function 39.
ompared to traditional sequence-based features, such as Amino Acid Composition (AAC) and Dipeptide Composition (DPC) 5, which mainly capture the frequency of individual amino acids or amino acid pairs, our spectral features incorporate additional structural information, including the adjacency between amino acid pairs, their positional distribution, and clustering behavior, offering a more expressive sequence characterization.
@@
In this section, we present our topology-enhanced machine learning model (Top-ML) and its results for ACPs classification. The natural vector, Magnus vector, terminus composition features, and peptide spectral representations are combined to form the input features for our Top-ML model. Figure 2 shows the general pipeline of our Top-ML model.
Model performance on AntiCP 2.0
@@
We train and assess the performance of our Top-ML model using the AntiCP 2.0 benchmark datasets 5. The AntiCP 2.0 datasets comprise Dataset A and Dataset B. Dataset A consists of 970 experimentally validated ACPs and 970 randomly selected peptides, assumed to be non-ACPs. Dataset B contains 861 experimentally validated ACPs and 861 non-ACPs, which are antimicrobial peptides (AMPs). AntiCP 2.0 (Datasets A and B) was partitioned into training and test sets using an 80/20 split. We follow the same protocol as previous studies: model parameters are tuned using 5-fold cross-validation on the training data, and the performance of the Top-ML model is evaluated on the test data using various performance metrics, including accuracy, sensitivity (true positive rate), specificity (true negative rate), and the Matthews correlation coefficient (MCC). The model selection process is detailed in the Model Selection. The final Top-ML model for AntiCP 2.0 is an Extra Trees classifier with 400 trees, each using a random subset of features equal to the square root of the total number of features to determine the optimal split. In this finalized model, peptide spectral representations are derived from the mean index position, and Magnus vectors are computed using a window size of 5. The terminus used to build terminal composition features is C15 for AntiCP 2.0 (Dataset A) and N15C15 for AntiCP 2.0 (Dataset B). The final Top-ML model has been further optimized by finding the optimal probability threshold that returns the best accuracy. All reported performances of the Top-ML model are medians over 100 repeated tests.
@@
Figure 2: Pipeline of topology-enhanced machine learning model (Top-ML) for ACPs prediction. (A) The peptide sequences from AntiCP 2.0 (Dataset A or B). (B) The peptide spectral features are obtained from the sequence-based Laplacian matrices. (C) The Magnus vectors, natural vectors, and terminal composition features are generated from peptide sequences. (D) The combined features serve as inputs for the classification of ACPs and non-ACPs.
@@
Table 1 summarizes the results of Top-ML and existing state-of-the-art machine learning models on AntiCP 2.0 (Datasets A and B). It can be observed that on AntiCP 2.0 (Dataset A), our Top-ML model achieves one of the highest accuracies and MCC scores, outperforming several existing state-of-the-art machine learning models. This is noteworthy because these models were built on deep learning algorithms while only tree-based classifiers have been used for Top-ML. Compared to AntiCP 2.0 5, which serves as a key benchmark model for comparison, our Top-ML model demonstrates an improvement of 1.3% in accuracy and 3.1% in specificity. On AntiCP 2.0 (Dataset B), the accuracy of Top-ML stands at 78.89%, which is higher than all other models compared, except for iACP-FSCM (82.50% accuracy) and ME-ACP(79.2%). When compared with the benchmark model AntiCP 2.0, there is an improvement of 3.5% in accuracy and 8.5% in specificity. Although iACP-FSCM and ME-ACP report higher accuracy, our Top-ML model can obtain a more balanced sensitivity and specificity score.